OpenStudy (anonymous):

Divide: (7 + 4i) ÷ (2 + 3i)

OpenStudy (espex):

Convert them to polar so you can divide the numbers and subtract the angles and then back for your final answer. :)

OpenStudy (espex):

Although you are probably using the conjugate method

OpenStudy (luigi0210):

Multiply by the conjugatge

OpenStudy (compassionate):

\[\frac{ 7 + 4i }{ 2 + 3i}\] Cancel like terms.

OpenStudy (anonymous):

another approach is to multiply by the denominators conjugate. \[\frac{ 7 + 4i }{ 2 + 3i } = \frac{ 7 + 4i }{ 2 + 3i } \times \frac{ 2- 3i }{ 2 - 3i } = \frac{ 14 - 13i - 12i^{2} }{ 4 - 9i^{2} }\]

OpenStudy (anonymous):

you still need to simplify :) remember that i^2 = -1.

OpenStudy (anonymous):

ans is 6i+26/13